A gas is placed in a closed container at 27 degree celcius if the temperature is raised to 327 degree celcius then pressure is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
As the container is closed,
hence volume is constant
p1 / T1 = p2 / T2
p1 /300  = p2 / 600
p2 = 2 * p1
hence pressure becomes double
 
Regards
Arun (askIITians forum expert)